OpenStudy (anonymous):

Use the Laplace transform to solve y"-2y'+2y=cos(t); y(0)=1, y'(0)=0.

OpenStudy (anonymous):

\[y''-2y'+2y=\cos t\] Applying the transform on both sides, you have \[\left(s^2Y(s)-sy(0)-y'(0)\right)-2\left(sY(s)-y(0)\right)+2Y(s)=\frac{s}{s^2+1}\\ Y(s)\left(s^2-2s+2\right)-s-0+2(1)=\frac{s}{s^2+1}\\ Y(s)\left(s^2-2s+2\right)=\frac{s}{s^2+1}+s-2\\ Y(s)\left(s^2-2s+2\right)=\frac{s+(s-2)(s^2+1)}{s^2+1}\\ Y(s)=\frac{s+s^3-2s^2+s-2}{\left(s^2+1\right)\left(s^2-2s+2\right)}\\ Y(s)=\frac{s^3-2s^2+2s-2}{\left(s^2+1\right)\left(s^2-2s+2\right)}\] Is this what you have so far? Just checking to see if I made any errors... Anyway, partial fraction decomposition should be the next step.

OpenStudy (anonymous):

How did you get the laplace transform for cos(t)? Is that also on the laplace transform table?

OpenStudy (anonymous):

It should be, yes. \[\mathcal{L}\left\{\cos at\right\}=\frac{s}{s^2+a^2}\] You could derive it from the definition of the transform.

OpenStudy (anonymous):

Also, do you suppose to factor out the numerator for Y(s)?

OpenStudy (anonymous):

Yes, that was the step between line one and two.

OpenStudy (anonymous):

What's the fastest way to factor out the long polynomial?

OpenStudy (anonymous):

Are you asking whether you can factor the numerator?

OpenStudy (anonymous):

Could you factor by grouping?

OpenStudy (anonymous):

No, I don't think so. You don't have to, PDF will work regardless.

OpenStudy (anonymous):

partial fraction decomp @Loser66

OpenStudy (anonymous):

@Idealist, do you know how to do that?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

But isn't it (s^3-2s^2+s-2)/((s^2+1)(s^2-2s+2))? What you got is (s^3-2s^2+2s-2)/((s^2+1)(s^2-2s+2)).

OpenStudy (anonymous):

\[\left(s^2Y(s)-sy(0)-y'(0)\right)-2\left(sY(s)-y(0)\right)+2Y(s)=\frac{s}{s^2+1}\\ s^2Y(s)-s-2sY(s)+2+2Y(s)=\frac{s}{s^2+1}\\ \left(s^2-2s+2\right)Y(s)=\frac{s}{s^2+1}+s-2\\ \left(s^2-2s+2\right)Y(s)=\frac{s}{s^2+1}+\frac{(s-2)(s^2+1)}{s^2+1}\\ \left(s^2-2s+2\right)Y(s)=\frac{s}{s^2+1}+\frac{s^3-2s^2+s-2}{s^2+1}\\ \left(s^2-2s+2\right)Y(s)=\frac{\color{red}s+s^3-2s^2+\color{red}s-2}{s^2+1} \]

OpenStudy (anonymous):

Never mind. I made a mistake, so sorry. Please continue.

OpenStudy (anonymous):

That's okay! \[Y(s)=\frac{s^3-2s^2+2s-2}{\left(s^2+1\right)\left(s^2-2s+2\right)}=\frac{As+B}{s^2+1}+\frac{Cs+D}{s^2-2s+2}\] \[s^3-2s^2+2s-2=(As+B)(s^2-2s+2)+(Cs+D)(s^2+1)\\ s^3-2s^2+2s-2=As^3+Bs^2-2As^2-2Bs+2As+2B+Cs^3+Ds^2+Cs+D\\ s^3-2s^2+2s-2=(A+C)s^3+(-2A+B+D)s^2+(2A-2B+C)s+(2B+D)\] Matching up coefficients yields the system \[\begin{cases}A+C=1\\-2A+B+D=-2\\2A-2B+C=2\\2B+D=-2 \end{cases}\]

OpenStudy (anonymous):

Solving that by hand looks tricky, so I'd suggest the use of a calculator if you can: http://www.wolframalpha.com/input/?i=a%2Bc%3D1%2C+-2a%2Bb%2Bd%3D-2%2C+2a-2b%2Bc%3D2%2C+2b%2Bd%3D-2

OpenStudy (anonymous):

On the test, were you allowed to solve that using a calculator?

OpenStudy (anonymous):

On a non-calculator test, you're likely to get a much easier problem where the coefficients are much easier to solve. If you're intent on solving this by hand, I'd suggest using a matrix method, where you would solve the matrix equation \[\begin{pmatrix}1&0&1&0\\-2&1&0&1\\2&-2&1&0\\0&2&0&1\end{pmatrix}\begin{pmatrix}A\\B\\C\\D\end{pmatrix}=\begin{pmatrix}1\\-2\\2\\-2\end{pmatrix}\]

OpenStudy (anonymous):

Okay, so once you get A, B, C, D, how do you write the Laplace?

OpenStudy (anonymous):

So you have \[Y(s)=\frac{1}{5}\left(\frac{s-2}{s^2+1}\right)+\frac{1}{5}\left(\frac{4s-6}{s^2-2s+2}\right)\] Take the inverse Laplace transform of both sides and you have \[y(t)=\frac{1}{5}\mathcal{L}^{-1}\left\{\frac{s-2}{s^2+1}\right\}+\frac{1}{5}\mathcal{L}^{-1}\left\{\frac{4s-6}{s^2-2s+2}\right\}\] Following so far? I think it would be easier to approach this one term at a time.

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

HAHAHA!

OpenStudy (anonymous):

Do you think you can take care of the inverse transforming? It shouldn't be too difficult if you break it up into multiple fractions.

OpenStudy (anonymous):

Yes. I get the idea.

OpenStudy (anonymous):

@Loser66, Of course there are plenty of tougher things than this. The only tough part about this problem is the algebra haha

OpenStudy (anonymous):

I know, right? I used to be good at algebra...

OpenStudy (anonymous):

I have question, what math subject comes after partial differential equations for physics and electrical engineering majors?

OpenStudy (anonymous):

I don't know about that, I'm just a math major. I get to have my pick of subjects.

OpenStudy (anonymous):

I see.